OpenStudy (anonymous):

Solve. Show work. -p-12=-4(8-p)

OpenStudy (anonymous):

Simplify the right-hand side: -p-12=-32+4p Add p to both sides: -12=-32+5p Add 32 to both sides: 20=5p Divide both sides by 5: 4=p.